Length of a minute hand of a clock is 4.5cm. find the avg velocity (in cm/s) of the tip of minute`s hand between:I) 6 AM to 6:30AM II) 6AM to 6:30 PM

Grade:11

2 Answers

Arun
6 years ago
Dear Preeti
 
Speed of hour hand=0.5 degrees per minute.
Speed of minutes hand=6 degrees per minute
angular distance moved by the minute hand between 

6.00 Am to 6.30 Am=30 (6 degrees ) =180 degrees
average angular velocity in this period is 
π/1800 rad/s
so, angular velocity of the tip of minute hand will be
4.5 x π/1800
=0.0025πcm/s
similarly for the second part the minute hand will cover( 6 

degrees)(750)=4500 degrees
Angular velocity in this period is
25π/750x60=π/1800 rad/s
Average velocity of the tip of minute hand will be
4.5 x π/1800=0.0025πcm/s
Biswajit
5 years ago
There minute hand is of length 4.5cm ,but for displacement we only need initial and final point,so minute hand tip is on 12 at 6am and then on 6 at 6.30 am so total displacement of its tip is diameter of clock I.e 2*4.5=9cm ,avg velocity is 9/(30*60) cm/s=5*10^-3cm/s.
2)that way you find between 6am to 6.30 pm I.e 9/(12.5*3600)=2*10^-4cm/s.